OpenStudy (haileyd):

What would be the graphing points of the system of equations? also please explain y=-2x-2 y=2x-2

OpenStudy (texaschic101):

you can make any points graphing points by subbing in a number for x and solving for y y = -2x - 2 ....lets make x = 1 y = -2(1) - 2 y = -2 - 2 y = -4 so when x = 1, y = -4....thus (1,-4) and do it like that till you have several points....for both lines. Now the solution for this system is when the two lines intersect. OR... you can find the x and y intercepts of each line. x int can be found by subbing in 0 for y and solving for x y int can be found by subbing in 0 for x and solving for y

OpenStudy (haileyd):

okay thank you, i understand it better now. i think i was just making it too complicated.